# Evaluating multiple models using labeled and unlabeled data

### v2

@@ D.1 Results reporting mean absolute error
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeling Dawid-Skene AutoEval Active-Testing SSME-KDE SSME-NF
AG News 20 1000 5.79 ± 3.04 5.88 ± 0.68 5.72 ± 4.16 8.31 ± 0.54 5.61 ± 2.77 8.60 ± 2.02 2.77 ± 0.96 5.56 ± 0.75
50 1000 4.09 ± 1.92 5.73 ± 0.87 2.97 ± 2.00 8.06 ± 0.68 3.68 ± 1.48 8.18 ± 2.28 2.72 ± 1.09 5.64 ± 1.03
100 1000 2.93 ± 1.52 5.52 ± 0.82 2.36 ± 1.48 7.66 ± 0.69 2.70 ± 1.29 8.23 ± 1.83 2.50 ± 1.09 5.32 ± 1.05
ImagenetBG 20 1000 6.62 ± 2.74 2.99 ± 0.90 33.45 ± 2.96 5.78 ± 0.71 6.55 ± 2.62 10.83 ± 5.34 8.76 ± 1.00 2.65 ± 0.67
50 1000 3.98 ± 1.63 3.01 ± 0.61 17.88 ± 2.78 5.69 ± 0.73 3.87 ± 1.56 12.25 ± 7.28 8.18 ± 0.90 2.66 ± 0.81
100 1000 2.97 ± 1.38 2.73 ± 0.57 9.37 ± 1.53 5.34 ± 0.63 2.73 ± 1.13 9.08 ± 4.22 8.02 ± 0.90 2.10 ± 0.68
MultiNLI 20 1000 7.46 ± 3.88 8.30 ± 0.81 7.95 ± 4.55 11.73 ± 0.55 7.20 ± 3.76 10.30 ± 4.14 1.98 ± 0.88 3.08 ± 0.65
50 1000 4.42 ± 1.99 8.14 ± 0.62 3.08 ± 2.25 11.41 ± 0.52 4.17 ± 1.96 11.85 ± 3.97 1.90 ± 0.76 2.79 ± 0.81
100 1000 3.27 ± 1.65 7.54 ± 0.69 2.47 ± 1.86 10.72 ± 0.54 3.17 ± 1.59 11.63 ± 4.14 2.02 ± 0.82 2.52 ± 0.77
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeling Dawid-Skene SSME-KDE SSME-NF
AG News 20 1000 7.04 ± 2.22 3.83 ± 0.35 4.48 ± 3.23 5.60 ± 0.28 2.24 ± 0.51 3.72 ± 0.50
50 1000 4.85 ± 1.54 3.75 ± 0.42 2.28 ± 1.36 5.37 ± 0.34 2.24 ± 0.59 3.81 ± 0.46
100 1000 3.24 ± 1.15 3.53 ± 0.37 1.89 ± 0.96 5.02 ± 0.40 2.15 ± 0.55 3.53 ± 0.60
ImagenetBG 20 1000 7.10 ± 2.79 2.73 ± 0.84 29.64 ± 2.84 4.76 ± 0.56 6.73 ± 0.54 2.49 ± 0.60
50 1000 4.00 ± 1.85 2.73 ± 0.52 14.18 ± 2.51 4.68 ± 0.48 6.42 ± 0.57 2.49 ± 0.72
100 1000 2.75 ± 1.13 2.56 ± 0.64 6.68 ± 1.04 4.54 ± 0.52 6.30 ± 0.62 1.96 ± 0.60
MultiNLI 20 1000 11.57 ± 4.06 3.87 ± 0.43 7.84 ± 4.12 2.95 ± 0.29 2.06 ± 0.88 1.75 ± 0.62
50 1000 6.14 ± 2.42 3.94 ± 0.32 3.10 ± 2.24 2.92 ± 0.37 2.06 ± 0.72 1.63 ± 0.59
100 1000 4.52 ± 1.83 3.82 ± 0.34 2.37 ± 1.66 3.18 ± 0.30 2.19 ± 0.76 1.45 ± 0.57
@@
D.2 Results reporting absolute performance estimates
[…]
Results thus far have reported aggregate errors in performance estimates across classifiers in the set. Here, we include results on a per-classifier basis in the context of toxicity detection, the task for which we have the largest variability in classifier quality (Tables S8, S9, S10, S11). The tables illustrate how SSME’s performance manifests on a per-classifier basis, often producing more accurate estimates than the baselines on the lowest performance classifiers. The tables also make evident that SSME’s improvement in performance estimation can be attributed to a significant reduction in the variance of performance estimates across different data splits.
@@ D.2 Results reporting absolute performance estimates
model Labeled Majority-Vote Pseudo-Labeled Dawid-Skene Bayesian-Calibration AutoEval Active-Testing SSME-KDE (Ours) Ground Truth
distilbert-CORAL 83.90 ± 7.16 87.68 ± 0.96 86.15 ± 4.11 84.75 ± 1.35 87.82 ± 4.41 85.04 ± 5.42 87.92 ± 10.82 86.49 ± 1.06 88.27 ± 0.09
distilbert-ERM 89.30 ± 6.15 97.63 ± 0.49 86.92 ± 3.35 95.08 ± 0.99 97.13 ± 0.96 90.79 ± 5.93 89.17 ± 8.24 93.59 ± 0.87 92.17 ± 0.07
distilbert-ERM-seed1 89.10 ± 7.12 97.48 ± 0.50 87.05 ± 3.34 94.86 ± 1.05 97.14 ± 1.01 90.75 ± 6.16 95.07 ± 6.47 93.54 ± 0.83 92.17 ± 0.08
distilbert-ERM-seed2 89.20 ± 6.34 98.10 ± 0.42 86.67 ± 3.43 95.78 ± 0.94 96.28 ± 1.23 90.43 ± 6.83 92.84 ± 6.25 93.65 ± 0.82 92.11 ± 0.08
distilbert-IRM 86.90 ± 7.62 93.03 ± 0.77 82.92 ± 3.26 95.27 ± 0.66 94.56 ± 2.34 88.11 ± 7.22 88.10 ± 9.13 91.65 ± 0.97 88.13 ± 0.09
distilbert-IRM-seed1 88.10 ± 7.35 94.03 ± 0.81 83.86 ± 3.31 95.92 ± 0.67 95.41 ± 1.76 89.27 ± 7.03 89.42 ± 9.13 92.26 ± 0.84 89.04 ± 0.10
distilbert-IRM-seed2 86.80 ± 6.91 93.49 ± 0.80 83.39 ± 3.24 95.55 ± 0.69 95.34 ± 2.18 87.78 ± 7.13 89.19 ± 9.26 92.08 ± 0.91 88.70 ± 0.08
model Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE (Ours) Ground Truth
distilbert-CORAL 13.80 ± 5.81 8.27 ± 0.79 8.78 ± 3.60 10.78 ± 1.08 7.47 ± 4.14 8.50 ± 0.95 5.98 ± 0.09
distilbert-ERM 10.37 ± 5.86 1.38 ± 0.44 11.33 ± 3.29 4.37 ± 1.06 1.97 ± 1.01 4.96 ± 0.75 6.14 ± 0.09
distilbert-ERM-seed1 9.91 ± 6.29 1.57 ± 0.38 11.29 ± 3.24 4.59 ± 1.13 1.93 ± 1.07 5.13 ± 0.78 6.21 ± 0.08
distilbert-ERM-seed2 10.02 ± 6.02 0.57 ± 0.32 10.96 ± 3.33 3.38 ± 0.99 2.32 ± 1.26 3.99 ± 0.71 4.94 ± 0.09
distilbert-IRM 12.59 ± 6.89 6.47 ± 0.70 15.77 ± 3.08 3.69 ± 0.67 4.61 ± 2.62 6.85 ± 0.88 10.61 ± 0.08
distilbert-IRM-seed1 11.93 ± 6.79 5.37 ± 0.73 15.03 ± 3.26 2.98 ± 0.63 3.88 ± 1.99 6.38 ± 0.93 9.78 ± 0.11
distilbert-IRM-seed2 12.06 ± 6.12 5.77 ± 0.75 15.55 ± 3.14 3.33 ± 0.74 3.87 ± 2.35 6.71 ± 0.89 10.18 ± 0.09
model Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE (Ours) Ground Truth
distilbert-CORAL 85.19 ± 14.15 95.09 ± 1.01 72.20 ± 6.89 94.89 ± 1.18 84.22 ± 3.49 91.38 ± 1.60 86.23 ± 0.17
distilbert-ERM 91.80 ± 7.11 99.29 ± 0.25 74.90 ± 7.69 98.64 ± 0.33 98.52 ± 0.94 95.97 ± 0.77 93.77 ± 0.11
distilbert-ERM-seed1 92.18 ± 7.46 99.22 ± 0.26 74.92 ± 7.68 98.58 ± 0.30 98.30 ± 0.95 95.96 ± 0.71 93.75 ± 0.10
distilbert-ERM-seed2 93.32 ± 5.93 99.47 ± 0.23 75.03 ± 7.74 98.89 ± 0.32 98.07 ± 1.12 96.16 ± 0.68 94.08 ± 0.10
distilbert-IRM 91.46 ± 8.74 98.41 ± 0.48 74.69 ± 7.49 98.28 ± 0.60 98.05 ± 0.96 95.38 ± 1.13 91.57 ± 0.13
distilbert-IRM-seed1 90.75 ± 10.48 98.20 ± 0.81 74.58 ± 7.58 97.97 ± 0.81 98.07 ± 1.02 95.18 ± 1.12 91.00 ± 0.16
distilbert-IRM-seed2 91.89 ± 8.13 98.40 ± 0.44 74.68 ± 7.60 98.39 ± 0.52 98.33 ± 0.94 95.59 ± 0.91 91.86 ± 0.11
model Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE (Ours) Ground Truth
distilbert-CORAL 60.78 ± 24.13 64.21 ± 5.46 31.91 ± 10.19 76.38 ± 5.00 50.86 ± 10.73 60.27 ± 5.50 40.00 ± 0.36
distilbert-ERM 77.25 ± 19.52 96.37 ± 1.09 42.28 ± 13.73 94.59 ± 1.22 92.51 ± 4.32 79.63 ± 3.68 72.19 ± 0.37
distilbert-ERM-seed1 79.38 ± 18.74 96.06 ± 1.18 42.27 ± 13.65 94.37 ± 1.15 91.54 ± 4.94 79.78 ± 3.46 72.30 ± 0.35
distilbert-ERM-seed2 79.11 ± 18.05 97.10 ± 1.01 42.47 ± 13.79 95.49 ± 1.19 90.15 ± 4.94 80.06 ± 3.34 73.33 ± 0.35
distilbert-IRM 77.74 ± 19.47 90.10 ± 2.39 40.79 ± 13.20 92.89 ± 1.47 88.23 ± 7.39 76.90 ± 4.22 65.86 ± 0.40
distilbert-IRM-seed1 79.21 ± 19.99 91.23 ± 2.66 41.23 ± 13.47 93.64 ± 1.62 89.58 ± 5.67 77.84 ± 3.91 66.50 ± 0.44
distilbert-IRM-seed2 77.63 ± 20.39 89.37 ± 2.40 40.65 ± 13.36 92.67 ± 1.77 90.34 ± 6.36 77.16 ± 3.67 65.46 ± 0.40
D.3 Comparison to baselines drawn from weak supervision
[…]

### v3

@@ D.3 Results reporting absolute performance estimates
model Labeled Majority-Vote PL Dawid-Skene SPE BC AutoEval Active-Testing SSME (Ours) Ground Truth
distilbert-CORAL 83.90 ± 14.04 83.81 ± 13.35 86.15 ± 8.05 84.75 ± 2.64 73.48 ± 51.60 87.82 ± 8.64 85.04 ± 10.63 87.92 ± 21.20 86.49 ± 2.07 88.27 ± 0.18
distilbert-ERM 89.30 ± 12.04 91.34 ± 7.53 86.92 ± 6.57 95.08 ± 1.94 84.55 ± 51.41 97.13 ± 1.88 90.79 ± 11.63 89.17 ± 16.14 93.59 ± 1.71 92.17 ± 0.14
distilbert-ERM-seed1 89.10 ± 13.96 91.28 ± 7.50 87.05 ± 6.54 94.86 ± 2.06 93.25 ± 10.36 97.14 ± 1.97 90.75 ± 12.08 95.07 ± 12.69 93.54 ± 1.63 92.17 ± 0.15
distilbert-ERM-seed2 89.20 ± 12.42 91.63 ± 7.32 86.67 ± 6.72 95.78 ± 1.84 85.49 ± 46.57 96.28 ± 2.42 90.43 ± 13.38 92.84 ± 12.24 93.65 ± 1.61 92.11 ± 0.15
distilbert-IRM 86.90 ± 14.93 92.67 ± 9.76 82.92 ± 6.39 95.27 ± 1.29 93.48 ± 14.96 94.56 ± 4.59 88.11 ± 14.15 88.10 ± 17.89 91.65 ± 1.91 88.13 ± 0.18
distilbert-IRM-seed1 88.10 ± 14.40 92.87 ± 8.98 83.86 ± 6.49 95.92 ± 1.32 95.76 ± 12.55 95.41 ± 3.44 89.27 ± 13.77 89.42 ± 17.90 92.26 ± 1.65 89.04 ± 0.19
distilbert-IRM-seed2 86.80 ± 13.54 92.48 ± 9.09 83.39 ± 6.35 95.55 ± 1.35 86.76 ± 52.55 95.34 ± 4.27 87.78 ± 13.98 89.19 ± 18.15 92.08 ± 1.78 88.70 ± 0.16
model Labeled Majority-Vote PL Dawid-Skene SPE BC SSME (Ours) Ground Truth
distilbert-CORAL 13.80 ± 11.38 12.60 ± 9.19 8.78 ± 7.05 10.78 ± 2.12 25.34 ± 53.26 7.47 ± 8.12 8.50 ± 1.87 5.98 ± 0.18
distilbert-ERM 10.37 ± 11.48 8.16 ± 7.77 11.33 ± 6.44 4.37 ± 2.08 14.87 ± 49.40 1.97 ± 1.97 4.96 ± 1.47 6.14 ± 0.17
distilbert-ERM-seed1 9.91 ± 12.33 8.34 ± 7.65 11.29 ± 6.35 4.59 ± 2.21 6.45 ± 10.72 1.93 ± 2.09 5.13 ± 1.53 6.21 ± 0.16
distilbert-ERM-seed2 10.02 ± 11.80 7.70 ± 7.75 10.96 ± 6.53 3.38 ± 1.93 14.36 ± 46.44 2.32 ± 2.47 3.99 ± 1.39 4.94 ± 0.19
distilbert-IRM 12.59 ± 13.50 6.52 ± 9.97 15.77 ± 6.04 3.69 ± 1.31 6.69 ± 14.92 4.61 ± 5.14 6.85 ± 1.72 10.61 ± 0.16
distilbert-IRM-seed1 11.93 ± 13.31 6.35 ± 9.18 15.03 ± 6.38 2.98 ± 1.24 4.18 ± 12.63 3.88 ± 3.89 6.38 ± 1.83 9.78 ± 0.21
distilbert-IRM-seed2 12.06 ± 12.00 6.71 ± 9.42 15.55 ± 6.16 3.33 ± 1.46 13.01 ± 51.77 3.87 ± 4.60 6.71 ± 1.74 10.18 ± 0.18
model Labeled Majority-Vote PL Dawid-Skene SPE BC SSME (Ours) Ground Truth
distilbert-CORAL 85.19 ± 27.74 90.53 ± 8.91 72.20 ± 13.51 94.89 ± 2.31 79.61 ± 72.95 84.22 ± 6.84 91.38 ± 3.14 86.23 ± 0.33
distilbert-ERM 91.80 ± 13.93 96.36 ± 8.70 74.90 ± 15.08 98.64 ± 0.65 90.62 ± 57.77 98.52 ± 1.84 95.97 ± 1.50 93.77 ± 0.22
distilbert-ERM-seed1 92.18 ± 14.63 96.36 ± 8.76 74.92 ± 15.05 98.58 ± 0.59 99.70 ± 0.58 98.30 ± 1.86 95.96 ± 1.39 93.75 ± 0.20
distilbert-ERM-seed2 93.32 ± 11.62 96.55 ± 8.86 75.03 ± 15.17 98.89 ± 0.62 90.81 ± 57.23 98.07 ± 2.20 96.16 ± 1.33 94.08 ± 0.19
distilbert-IRM 91.46 ± 17.13 95.52 ± 13.36 74.69 ± 14.67 98.28 ± 1.17 99.06 ± 3.55 98.05 ± 1.88 95.38 ± 2.22 91.57 ± 0.25
distilbert-IRM-seed1 90.75 ± 20.55 95.25 ± 12.97 74.58 ± 14.85 97.97 ± 1.60 99.52 ± 1.60 98.07 ± 2.00 95.18 ± 2.19 91.00 ± 0.31
distilbert-IRM-seed2 91.89 ± 15.94 95.37 ± 13.62 74.68 ± 14.90 98.39 ± 1.02 90.44 ± 57.04 98.33 ± 1.84 95.59 ± 1.79 91.86 ± 0.23
model Labeled Majority-Vote PL Dawid-Skene SPE BC SSME (Ours) Ground Truth
distilbert-CORAL 60.78 ± 47.30 67.66 ± 48.12 31.91 ± 19.96 76.38 ± 9.81 71.94 ± 75.73 50.86 ± 21.03 60.27 ± 10.78 40.00 ± 0.70
distilbert-ERM 77.25 ± 38.26 82.78 ± 55.03 42.28 ± 26.91 94.59 ± 2.39 79.50 ± 71.95 92.51 ± 8.47 79.63 ± 7.22 72.19 ± 0.73
distilbert-ERM-seed1 79.38 ± 36.73 83.34 ± 53.48 42.27 ± 26.76 94.37 ± 2.25 83.36 ± 69.87 91.54 ± 9.68 79.78 ± 6.78 72.30 ± 0.68
distilbert-ERM-seed2 79.11 ± 35.38 83.41 ± 55.47 42.47 ± 27.03 95.49 ± 2.34 78.44 ± 73.38 90.15 ± 9.69 80.06 ± 6.54 73.33 ± 0.69
distilbert-IRM 77.74 ± 38.16 81.26 ± 57.61 40.79 ± 25.87 92.89 ± 2.88 81.71 ± 76.90 88.23 ± 14.48 76.90 ± 8.27 65.86 ± 0.78
distilbert-IRM-seed1 79.21 ± 39.17 81.73 ± 57.61 41.23 ± 26.40 93.64 ± 3.17 82.29 ± 74.37 89.58 ± 11.12 77.84 ± 7.67 66.50 ± 0.86
distilbert-IRM-seed2 77.63 ± 39.97 80.53 ± 57.24 40.65 ± 26.19 92.67 ± 3.46 77.63 ± 77.91 90.34 ± 12.47 77.16 ± 7.20 65.46 ± 0.79
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene AutoEval Active-Testing SSME (Ours)
AG News 20 1000 5.48 ± 0.89 5.13 ± 0.16 11.72 ± 1.23 6.70 ± 0.13 5.59 ± 1.18 8.09 ± 0.80 3.72 ± 0.19
50 1000 3.92 ± 0.63 4.96 ± 0.14 5.22 ± 0.78 6.51 ± 0.12 4.36 ± 0.78 8.88 ± 1.40 3.54 ± 0.17
100 1000 2.71 ± 0.45 4.74 ± 0.19 2.69 ± 0.48 6.20 ± 0.16 2.71 ± 0.59 8.77 ± 1.20 3.41 ± 0.24
MultiNLI 20 1000 7.46 ± 1.07 8.30 ± 0.35 7.95 ± 1.26 11.73 ± 0.15 7.20 ± 1.04 10.30 ± 1.77 1.98 ± 0.24
50 1000 4.42 ± 0.55 8.14 ± 0.27 3.08 ± 0.62 11.41 ± 0.14 4.17 ± 0.54 11.85 ± 1.70 1.90 ± 0.21
100 1000 3.27 ± 0.46 7.54 ± 0.30 2.47 ± 0.51 10.72 ± 0.15 3.17 ± 0.44 11.63 ± 1.77 2.02 ± 0.23
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene SSME (Ours)
AG News 20 1000 5.53 ± 0.93 4.30 ± 0.13 11.85 ± 1.21 5.80 ± 0.11 3.37 ± 0.17
50 1000 3.88 ± 0.76 4.18 ± 0.11 5.36 ± 0.77 5.70 ± 0.10 3.28 ± 0.15
100 1000 2.61 ± 0.49 4.07 ± 0.15 2.78 ± 0.48 5.53 ± 0.13 3.13 ± 0.21
MultiNLI 20 1000 11.57 ± 1.13 3.87 ± 0.18 7.84 ± 1.14 2.95 ± 0.08 2.06 ± 0.24
50 1000 6.14 ± 0.67 3.94 ± 0.14 3.10 ± 0.62 2.92 ± 0.10 2.06 ± 0.20
100 1000 4.52 ± 0.51 3.82 ± 0.15 2.37 ± 0.46 3.18 ± 0.08 2.19 ± 0.21
@@
model SSME-KDE-gaussian SSME-KDE-epanechnikov SSME-KDE-exponential
distilbert_CORAL 86.61 ± 1.3587 86.61 ± 1.3586 86.60 ± 1.3884
distilbert_ERM 93.83 ± 1.0066 93.83 ± 1.0067 93.81 ± 0.9879
distilbert_ERM_seed1 93.76 ± 0.9585 93.75 ± 0.9584 93.73 ± 0.9314
distilbert_ERM_seed2 93.77 ± 0.6859 93.77 ± 0.6859 93.75 ± 0.6643
distilbert_IRM 91.80 ± 1.0459 91.80 ± 1.0460 91.78 ± 1.0122
distilbert_IRM_seed1 92.37 ± 1.0373 92.37 ± 1.0373 92.35 ± 1.0252
distilbert_IRM_seed2 92.33 ± 0.9593 92.33 ± 0.9592 92.31 ± 0.9346
Table S13: Robustness of SSME to kernel choice. Because density estimation can be sensitive to kernel choice, we analyze the stability of SSME’s performance estimates across three kernel types. We find no significant imapct of kernel choice on SSME’s performance estimates (although different kernels do produce small changes in accuracy estimates). Results hold across metrics and tasks; we report results for each toxicity detection classifier and accuracy for brevity.
D.4 Results reporting robustness to kernel choice
We find that SSME’s performance estimates are stable across three kernel choices: a Gaussian kernel, an Epachenikov kernel, and an exponential kernel. Table S13 reports our results in the context of toxicity detection and accuracy, but results generalize across tasks and metrics.
D.5 Results reporting performance of alternate parameterization
SSME can be fit using other functions to approximate $P(\mathbf{s}|y)$. We additionally explored the value of parameterizing each component using a normalizing flow (implementation details in Sec. E.3), a technique that is increasingly used to approximate complex densities [2].
We find that using a KDE outperforms the normalizing flow on each binary task, for every metric (except estimating AUPRC for SARS-CoV inhibition prediction, Table S15). We find that the KDE performs similarly to the normalizing flow for each LLM-based task, although by a smaller margin (Table S14). We hypothesize that this is due to the established challenges of approximating multi-modal distributions with normalizing flows [65, 16].
For higher-dimensional tasks, the normalizing flow seems to outperform the KDE. In particular, we ran an experiment on ImagenetBG, which contains nine classes and four classifiers. Table S16 demonstrates that in this setting, the normalizing flow performs better than the KDE.
@@ D.5 Results reporting performance of alternate parameterization
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene AutoEval Active-Testing SSME (Ours) SSME-NF
AG News 20 1000 5.48 ± 0.45 5.13 ± 0.08 11.72 ± 0.63 6.70 ± 0.07 5.59 ± 0.60 8.09 ± 0.41 3.72 ± 0.10 3.52 ± 0.11
50 1000 3.92 ± 0.32 4.96 ± 0.07 5.22 ± 0.40 6.51 ± 0.06 4.36 ± 0.40 8.88 ± 0.72 3.54 ± 0.09 3.50 ± 0.11
100 1000 2.71 ± 0.23 4.74 ± 0.10 2.69 ± 0.25 6.20 ± 0.08 2.71 ± 0.30 8.77 ± 0.61 3.41 ± 0.12 3.40 ± 0.13
MultiNLI 20 1000 7.46 ± 0.55 8.30 ± 0.18 7.95 ± 0.64 11.73 ± 0.08 7.20 ± 0.53 10.30 ± 0.90 1.98 ± 0.12 3.08 ± 0.09
50 1000 4.42 ± 0.28 8.14 ± 0.14 3.08 ± 0.32 11.41 ± 0.07 4.17 ± 0.28 11.85 ± 0.87 1.90 ± 0.11 2.79 ± 0.11
100 1000 3.27 ± 0.23 7.54 ± 0.15 2.47 ± 0.26 10.72 ± 0.08 3.17 ± 0.23 11.63 ± 0.90 2.02 ± 0.12 2.52 ± 0.11
Table S14: Mean absolute error in accuracy estimation on multiclass tasks, including NF parametrization. SSME-NF performs comparably or worse that SSME (Ours) across the two multiclass tasks and three labeled data settings we consider.
Dataset $n_{\ell}$ $n_{u}$ Labeled SSME-M SSME (Ours) SSME-NF
Critical Outcome 20 1000 5.19 ± 1.07 1.70 ± 0.27 0.67 ± 0.13 9.80 ± 1.41
50 1000 2.90 ± 0.59 1.65 ± 0.25 0.78 ± 0.13 8.81 ± 1.11
100 1000 2.09 ± 0.41 1.30 ± 0.20 0.77 ± 0.13 8.28 ± 1.64
ED Revisit 20 1000 5.11 ± 0.98 1.64 ± 0.34 0.45 ± 0.10 3.54 ± 1.87
50 1000 2.02 ± 0.58 1.46 ± 0.27 0.53 ± 0.11 2.63 ± 1.10
100 1000 1.43 ± 0.32 1.18 ± 0.25 0.57 ± 0.11 2.51 ± 0.94
Hospital Admission 20 1000 7.32 ± 1.25 3.29 ± 0.47 1.88 ± 0.29 3.19 ± 0.95
50 1000 5.40 ± 0.83 3.17 ± 0.51 1.95 ± 0.27 2.27 ± 0.62
100 1000 3.64 ± 0.55 3.06 ± 0.45 1.51 ± 0.23 3.66 ± 2.31
SARS-CoV Inhibition 20 1000 6.11 ± 0.96 3.06 ± 0.23 2.30 ± 0.16 11.67 ± 5.26
50 1000 3.22 ± 0.57 2.59 ± 0.26 2.35 ± 0.10 20.89 ± 7.38
100 1000 2.04 ± 0.38 1.84 ± 0.24 2.36 ± 0.13 9.89 ± 4.52
Toxicity Detection 20 1000 5.95 ± 0.73 6.71 ± 0.23 2.34 ± 0.15 3.44 ± 0.11
50 1000 4.03 ± 0.68 5.38 ± 0.28 2.22 ± 0.13 3.31 ± 0.13
100 1000 2.43 ± 0.41 3.80 ± 0.32 2.14 ± 0.15 3.32 ± 0.13
Table S15: Mean absolute error in accuracy estimation on binary tasks, including NF parameterization. We include columns for the Labeled, SSME fit to each classifier individually (SSME-M), SSME parameterized by a KDE (SSME (Ours)), and SSME parameterized by an NF (SSME-NF). While SSME-NF sometimes outperforms SSME-M, SSME-NF never outperforms SSME-KDE on accuracy estimation in any of the tasks or amounts of labeled data we consider.
Dataset $\mathbf{n_{\ell}}$ $\mathbf{n_{u}}$ Labeled Majority-Vote PL DS AutoEval Active-Testing SSME (KDE) SSME (NF)
ImnetBG 20 1000 6.62 $\pm$ 2.74 2.99 $\pm$ 0.90 33.45 $\pm$ 2.96 5.78 $\pm$ 0.71 6.55 $\pm$ 2.62 10.83 $\pm$ 5.34 8.76 $\pm$ 1.00 2.65 $\pm$ 0.67
ImnetBG 50 1000 3.98 $\pm$ 1.63 3.01 $\pm$ 0.61 17.88 $\pm$ 2.78 5.69 $\pm$ 0.73 3.87 $\pm$ 1.56 12.25 $\pm$ 7.28 8.18 $\pm$ 0.90 2.66 $\pm$ 0.81
ImnetBG 100 1000 2.97 $\pm$ 1.38 2.73 $\pm$ 0.57 9.37 $\pm$ 1.53 5.34 $\pm$ 0.63 2.73 $\pm$ 1.13 9.08 $\pm$ 4.22 8.02 $\pm$ 0.90 2.10 $\pm$ 0.68
Table S16: Comparison across evaluation methods on the ImnetBG dataset with varying numbers of labeled examples ($n_{\ell}$) and unlabeled examples ($n_{u}$). SSME (NF) consistently achieves the lowest MAE.
D.6 Results reporting robustness to initializations
We find that SSME’s performance remains strong when using an alternative plausible initialization method: using a majority vote among k=5 nearest neighbors among the labeled examples to initialize cluster assignment for unlabeled examples. We compare this alternate initialization method to our original initialization method (which we refer to as “draw" in the table) on the CivilComments dataset with three different amounts of labeled data (20, 50, and 100 points). We include our results on ECE and accuracy estimation error in the table S17, where we find that the two initialization methods perform comparably on the CivilComments dataset.
Labeled Examples Initialization MAE, ECE MAE, Accuracy
20 KNN 0.045 0.046
20 Draw 0.035 0.034
50 KNN 0.037 0.040
50 Draw 0.043 0.043
100 KNN 0.031 0.033
100 Draw 0.037 0.036
Table S17: Performance comparison across different numbers of labeled examples and initialization methods (Dataset: CivilComments).
D.7 Results reporting performance as number of classifiers increases
As shown in table S18, we find that SSME’s performance improves as the number of classifiers increases, which accords with our theoretical results. As a caveat, our theory predicts that SSME is likely not to perform well if one adds an inaccurate (e.g., worse than random) classifier to the set.
We also find that SSME’s performance does not vary greatly across classifier sets, suggesting that SSME’s performance is robust to using different classifier sets. In particular, the standard deviation in accuracy estimation error across random classifier sets is always under 0.011, and the standard deviation in ECE estimation error is always under 0.015, indicating consistent performance across classifier sets.
First, we find that performance does not vary greatly across classifier sets, suggesting that SSME’s performance is robust to using different classifiers. In particular, the standard deviation in accuracy estimation error across random classifier sets is always under 0.011, and the standard deviation in ECE estimation error is always under 0.015.
Second, we find that SSME’s performance improves as the number of classifiers increases, which accords with our theoretical results. As a caveat, our theory predicts that SSME is likely not to perform well if one adds an inaccurate (e.g., worse than random) classifier to the set.
# Classifiers MAE, ECE (SD) MAE, Accuracy (SD)
2 0.0457 (0.0059) 0.0471 (0.0047)
3 0.0302 (0.0107) 0.0308 (0.0098)
4 0.0268 (0.0094) 0.0273 (0.0096)
5 0.0252 (0.0072) 0.0253 (0.0078)
6 0.0235 (0.0043) 0.0235 (0.0048)
7 0.0225 (only one subset) 0.0223 (only one subset)
Table S18: Performance as the number of classifiers increases. Standard deviations (SD) are shown in parentheses. Dataset: CivilComments.
D.8 Results reporting performance for different bandwith selection procedures
We ran an experiment to test the robustness of our results to the choice of bandwidth. In the paper, we use the Sheather-Jones algorithm to automatically identify a bandwidth. We compare this approach to (1) another automated bandwidth selection algorithm (the Silverman algorithm) and (2) two fixed bandwidths, chosen to be larger but within an order of magnitude of the Sheather-Jones selected bandwidth. We conduct experiments using the CivilComments dataset and 3 labeled dataset sizes (20, 50, and 100 labeled datapoints). While results remain strong in all cases, we achieve the best performance when using the Sheather-Jones bandwidth selection procedure.
# Labeled Examples Bandwidth Selection Procedure MAE, ECE MAE, Accuracy
20 Sheather–Jones 0.024 0.023
20 Silverman 0.040 0.038
20 1.0 0.037 0.036
20 2.0 0.046 0.044
50 Sheather–Jones 0.023 0.022
50 Silverman 0.066 0.065
50 1.0 0.062 0.062
50 2.0 0.045 0.044
100 Sheather–Jones 0.022 0.021
100 Silverman 0.049 0.047
100 1.0 0.056 0.055
100 2.0 0.034 0.032
Table S19: (Dataset: CivilComments) Comparison of bandwidth selection procedures. Sheather–Jones outperforms another automated selection method (Silverman) and two fixed bandwidths.
D.9 Comparison to baselines drawn from weak supervision